Alicia is arranging photographs of five family members in a row on her refrigerator. If she wants the photograph of her mother and the photograph of her father to be on opposite ends of the row, how many arrangements for the photograph are possible?
Answer by ikleyn(52781)

.

There are 3! = 1*2*3 permutations (arrangements) for three other members of the family (excluding photos of the parents),

and 2 arrangements for the parents' photos.


So, in all, there are  2*3! = 2*6 = 12  different arrangements.

Solved.